# Over-the-Air DPD and Reciprocity Calibration in Massive MIMO and Beyond

### v2

@@ III-C Reciprocity calibration
The last stage consists of performing OTA reciprocity calibration considering the TX-chains previously linearized through the DPD linearization stage. To this end, each BS antenna transmits inter-antenna pilots to other antennas. The received symbols at the $j$’th antenna from the $i$’th antenna may be expressed as
@@
In [6], it was noted that multiplying all the reciprocity parameters by a common scalar does not compromise the effectiveness of the reciprocity calibration. [^5] Note that this constant may be absorbed in the linear response of the UEs RXs-chain, given by $\boldsymbol{R}_{\mathrm{U}}$ in (3). Thus, we may select one of the calibration parameters, e.g., $c_{1}$, and normalize all the rest by that value. The resulting scaled calibration parameters may then be expressed as
@@
The estimation error of $\widehat{\tilde{c}_{i}}$ can be reduced by averaging several estimates of $r_{1}\tilde{t}_{i}$ and $r_{i}\tilde{t}_{1}$, which is possible if each BS antenna transmits $N_{\text{cal}}\geq 2$ pilots in (19). Note further that, assuming reciprocity of the mutual coupling coefficients, i.e. $h_{ij}=h_{ji}$, we could still estimate $\tilde{c}_{m}$ using (19) without explicit knowledge of $h_{ij}$ since the coefficients would cancel each other in (22).
We have thus shown that we can estimate all the entries of the calibration matrix up to a constant, i.e., we can estimate $\tilde{\boldsymbol{C}}=\frac{1}{c_{1}}\boldsymbol{C}$, by means of simple linear estimators. This allows us to achieve reciprocity calibration without the need for high complexity iterative algorithms, such as the algorithms used in [6].
@@ IV Numerical results
In this section, we perform simulations to validate the feasibility and assess the performance of the proposed OTA-based DPD and calibration method. The number of BS antennas and the number of single-antenna UEs are $M=100$ and $K=10$, respectively. For the BS TX-chains non-linearity parameters in (5), we fit a 3rd order polynomial to the measurement data from [13] for a Gallium Nitride (GaN) amplifier operating at 2.1 GHz at a sample rate of 200 MHz and a signal bandwidth of 40 MHz. For the RX-chains complex gains, we use the values in [6] given by $r_{m}^{B}=0.9+0.2\frac{M-m}{M}\exp(j2\pi m/M)$. To perform the DPD, we generate a look-up table based on the OTA data to implement the inverse function of the non-linear response. For the mutual coupling channel gains in (9), we have used the linear LS fit based on the measurements in [6]. We perform the simulations for different levels of a reference OTA link SNR, given by the RX-SNR for the link between the antennas with least mutual coupling gain.
@@
Fig. 1 illustrates the average mean square error (MSE) of the calibration matrix estimation with our proposed OTA-based method for different levels of reference SNR. The calibration matrix is estimated after performing the OTA non-linearity characterization, which is performed with $N_{\text{dpd}}=500$ or $N_{\text{dpd}}=2000$ OTA transmissions. In the calibration step, we have considered $N_{\text{cal}}=200$ or $N_{\text{cal}}=500$ OTA transmissions. Note that transceiver characteristics are slowly-varying parameters [6], which means that large values of $N_{\text{dpd}}$ and $N_{\text{cal}}$ would still have a rather small impact on the total overhead. For comparison, we have also included the case with unlimited $N_{\text{dpd}}$, and the CRLB derived in [6], which assumes linear TX-chains. We can see that the calibration matrix estimation MSE for the proposed method is fairly close to the CRLB for [6], even though we are considering non-linear TX-chains. Note that, in order to approach the CRLB, which gives a theoretical bound on the achievable performance, [6] requires (at high SNR) an iterative algorithm of considerable complexity. Specifically, the reported complexity order in [6], which further outperforms previous state-of-the-art [5], is given by $\mathcal{O}(M^{2}N_{\text{ite}})$, where $N_{\text{ite}}$ is the number of iterations of the algorithm. On the other hand, in our reciprocity calibration method the complexity is given by $\mathcal{O}(MN_{\text{cal}})$, since it requires averaging $N_{\text{cal}}$ numbers where each of them is obtained by performing 2 multiplications for each of the $M-1$ antennas (excluding the reference one). Thus, for massive MIMO and beyond, where $M\gg 1$, our method may attain a significant complexity reduction. Moreover, the effect of reducing $N_{\text{cal}}$ only has a minor effect on the calibration matrix MSE, as seen from Fig. 1. We can also see that the MSE with the proposed method can approach the perfect OTA-DPD case for given $N_{\text{cal}}$, which corresponds to infinite $N_{\text{dpd}}$ in (14).
@@
Figure 1: Average MSE of calibration matrix estimation using the proposed OTA-based DPD and Calibration.
@@
Fig. 2 illustrates the CDF of DL data rate under ZF precoding, where we have generated $10^{4}$ DL channel realizations with i.i.d. Rayleigh fading channel, and the DL signal power for each UE is selected to achieve an average SNR of $10$ dB at their receivers. We have selected $N_{\text{cal}}=500$ for the OTA reciprocity calibration. For comparison, we have included two extreme cases with perfect DPD as benchmarks, one ideal case with perfect DL CSI available, and one baseline case with no calibration. To have a fair comparison with the benchmarks in terms of DPD performance, and since the results in Fig. 1 show very close MSE performance between limited-size and infinite-size OTA-DPD cases, we consider the infinite-size OTA-DPD scenario for the OTA plots. We can see that, for moderate values of the reference SNR, our proposed method can significantly improve the DL data rate, and it can approach the ideal case with perfect DL CSI without requiring high-complexity iterative algorithms and dedicated hardware for performing DPD and reciprocity calibration.
@@ V Conclusion
In this paper, we have proposed an OTA-based method for DPD and reciprocity calibration in massive MIMO systems and beyond. In particular, we considered a memory-less non-linearity model for the BS transmitters and proposed to linearize the transmitters and perform the calibration by using OTA measurements of the mutual coupling among the BS antennas. We showed that, by only using the OTA data, we can effectively linearize the transmitters and perform reciprocity calibration with reduced complexity over state-of-the-art. Simulation results showed promising performance of the proposed methodology, both in terms of the calibration matrix estimation error and DL data-rates when applying ZF precoding after using our OTA-based DPD and calibration method.
Figure 2: CDF of UEs DL Data Rate.

### v3

@@ III-C Reciprocity calibration
The last stage consists of performing OTA-based calibration considering the TX-chains previously linearized through the OTA-DPD stages. To this end, each BS antenna transmits pilots to other antennas. The received symbols at the $j$’th antenna from the $i$’th antenna may be expressed as
@@
In [5], it was noted that multiplying all the reciprocity parameters by a common scalar does not compromise the effectiveness of the reciprocity calibration. [^5] This constant scalar may be absorbed in the linear response of the UEs RX-chains, given by $\boldsymbol{R}_{\mathrm{U}}$ in (3). Thus, we may select one of the calibration parameters, e.g., $c_{1}$, and normalize all the rest by that value. The resulting scaled calibration parameters may then be expressed as
@@
The estimation error of $\widehat{\tilde{c}_{i}}$ can be reduced by averaging several estimates of $r_{1}\tilde{t}_{i}$ and $r_{i}\tilde{t}_{1}$, which is possible if each BS antenna transmits $N_{\text{cal}}\geq 2$ pilots in (19). Note further that, assuming reciprocity of the mutual coupling coefficients, i.e. $h_{ij}=h_{ji}$, we could still estimate $\tilde{c}_{m}$ using (19) without explicit knowledge of $h_{ij}$ since the coefficients would cancel each other in (22). The method does not rely on this assumption, since the coupling coefficients can be estimated in practice. Nevertheless, the coupling coefficients between antennas are reciprocal per definition since we are absorbing the non-reciprocity within our model.
We have thus shown that we can estimate all the entries of the calibration matrix up to a constant, i.e., we can estimate $\tilde{\boldsymbol{C}}=\frac{1}{c_{1}}\boldsymbol{C}$, by means of simple linear estimators. This allows achieving reciprocity without the need for high complexity iterative algorithms, such as the algorithms used in [5].
@@ IV Numerical results
In this section, we perform simulations to validate the feasibility and assess the performance of the proposed method. The number of BS antennas and the number of single-antenna UEs are $M=100$ and $K=10$, respectively. For the BS TX-chains non-linearity parameters in (5), we fit a 3rd order polynomial to the measurement data from [14] for a Gallium Nitride (GaN) amplifier operating at 2.1 GHz at a sample rate of 200 MHz and a signal bandwidth of 40 MHz. For the RX-chains complex gains, we use the values in [5] given by $r_{m}^{B}=0.9+0.2\frac{M-m}{M}\exp(j2\pi m/M)$. To implement the imperfect inverse function in the DPD, we generate a look-up table based on the OTA data. For the mutual coupling channel gains in (9), we have used the linear LS fit based on the measurements in [5]. We also define the OTA link signal to noise ratio (SNR) as the receive SNR for the link between the antennas with least mutual coupling gain.
@@
Fig. 1 illustrates the average mean square error (MSE) of the calibration matrix estimation with our proposed OTA-based method for different levels of SNR. The calibration matrix is estimated after performing the OTA-DPD, with $N_{\text{dpd}}=500$ or $2000$ OTA transmissions. In the calibration step, we have considered $N_{\text{cal}}=200$, $500$, or $2000$ OTA transmissions. Note that transceiver characteristics are slowly-varying parameters, which means that even larger values of $N_{\text{dpd}}$ and $N_{\text{cal}}$ would still have a rather small impact on the total overhead. For comparison we have included an upper bound for the performance of the reciprocity calibration problem at hand, CRLB, derived in [5], which assumes linear TX-chains. We have also included the case with an ideal DPD followed by the OTA-based reciprocity calibration. Firstly, we can see that the performance of the OTA-based DPD and reciprocity calibration approaches the ideal DPD case. Secondly, we can see that the performance is fairly close to the CRLB, even though we have the extra challenge of dealing with non-linear TX-chains. Note that, in order to approach the CRLB, which is the ultimate performance bound for the reciprocity problem, [5] requires an iterative algorithm of considerable complexity. The reported complexity order in [5], which is considered state-of-the-art in massive MIMO reciprocity calibration [16], is given by $\mathcal{O}(M^{2}N_{\text{ite}})$, where $N_{\text{ite}}$ is the number of iterations of the algorithm. The complexity of our reciprocity calibration method is given by $\mathcal{O}(MN_{\text{cal}})$, since it requires averaging $N_{\text{cal}}$ numbers where each of them is obtained by performing 2 multiplications for each of the M1M-1 antennas. Thus, for massive MIMO and beyond, where $M\gg 1$, our method may even attain significant complexity reduction, since reducing $N_{\text{cal}}$ only has a minor impact on the calibration matrix MSE, as seen from Fig. 1. Other state-of-the-art reciprocity calibration methods require even higher complexity, while their performance is still limited by the CRLB. For example, the one in [7] for mm-Wave systems requires $\mathcal{O}(M^{3})$ complexity to tackle the non-reciprocity problem with linear transceivers.
@@
Figure 1: Average MSE of calibration matrix estimation using the proposed OTA DPD and reciprocity calibration.
@@
Fig. 2 illustrates the CDF of DL data rate under ZF precoding, where we have generated $10^{4}$ DL realizations of an i.i.d. Rayleigh fading channel, and the DL signal power for each UE is selected to achieve an average SNR of $10$ dB at their receivers. We have selected $N_{\text{cal}}=500$ for the OTA reciprocity calibration, OTA reference SNR of 0 dB, and OTA-DPD with $N_{\text{dpd}}=500$. We have also performed the same reciprocity calibration for a case with perfect DPD. For comparison, we have included two extreme cases both with a perfect DPD, one with perfect DL CSI (ideal calibration), and one worst-case scenario with no calibration. We can observe that the limited-size DPD performs very close to the perfect DPD case, which further confirms the observations from Fig. 1. As for the calibration performance, we can see that the proposed OTA-DPD and reciprocity calibration approaches the ideal case with perfect DL CSI, without requiring high-complexity iterative algorithms, even-though we are dealing with the extra challenge of non-linear TX-chains. The gain from adopting the proposed OTA-case is more significant for higher number of antennas, i.e., massive MIMO and beyond, and the OTA-DPD performs closer to the ideal DPD case.
@@ V Conclusion
In this paper, we have proposed an OTA-based method for DPD and reciprocity calibration in massive MIMO and beyond. In particular, we considered a memory-less non-linearity model for the BS transmitters and proposed to perform the linearization and reciprocity calibration by using OTA measurements of the mutual coupling among the BS antennas. We showed that, by only using the OTA data, we can effectively linearize the transmitters and perform reciprocity calibration with reduced complexity over state-of-the-art. Simulation results showed promising performance of the proposed methodology, both in terms of the calibration matrix estimation error and the DL data-rates when applying ZF precoding after our OTA-based DPD and calibration method.
Figure 2: CDF of a UE DL data rate.